The enthalpy of combustion of benzoic acid (C6H5COOH) is commonly used as the standard for calibrating constant-volume bomb calorimeters; its value has been accurately determined to be – 3226.7 kJ/mol. (a) When 0.4654 g of a-D-glucose (C6H12O6) was oxidized the temperature rose from 21.22 °C to 22.28 °C. Calculate the enthalpy of combustion of glucose and the value of DUrxn for the combustion. (Hint: Be sure to write a balanced equation for each reaction first and assume the product H2O is a liquid).
Answer #1

equation of the reaction is given by C6H12O6 (s) + 6O2(g) à 6CO2(g) + 6H2O(l)

q = C D T = nglu D cU o gives D cU o = (6.81 kJ/K) 1.06K (180 g/mol)/0.4654 g or D cU o = - 2792 kJ/mol

Note that D cU o + D (pV) = D cU o + D (nRT)o = D cH o (assuming a perfect gas). Since D n = net change in #moles of gas = 0, D cU o = D cH o.

Note D cH o = - 2808 kJ/mol in A1.1

Find D fH o for 6C(s) + 6H2(g) + 3O2(g) à C6H12O6 (s)

6CO2(g) + 6H2O(l) à C6H12O6 (s) + 6O2(g) D cH o = +2792 kJ/mol

6C(s) + 6O2(g) à 6CO2(g) 6 x D fH o = - 2361

6H2(g) + 3O2(g) à 6H2O(l) 6 x D fH o = -1715

Add these 3 eqns together to get the enthalpy of formation for glucose

D fH o = - 1283 kJ/mol Note D fH o = - 1274 kJ/mol in A1.1